# DPImageBench: A Unified Benchmark for Differentially Private Image Synthesis

### v2

@@ 3 Overview of DPImageBench
DPImageBench should comprehensively evaluate the quality of synthetic images generated by different algorithms for comparative studies.
@@
(1) Model Selection. DPImageBench allows users to select among twelve distinct DP image synthesis algorithms, which are categorized into three groups as introduced in Section 4. DPImageBench also provides simple interfaces to allow users to integrate new algorithms.
(2) Data Preparation. This module prepares the public images and sensitive images for the following modules by partitioning and preprocessing the data, such as normalization. DPImageBench offers a flexible tool that allows users to select public and sensitive datasets.
(3) Model Training. This module involves (optional) synthesizer pretraining on the public dataset and the main training on the sensitive dataset. We unify the model architectures and hyper-parameters across different methods. Besides, we provide flexible configuration files, allowing users to adjust hyper-parameters according to their preferences. DPImageBench supports pretraining for all algorithms, even those not originally designed with pretraining. Details of the pretraining methods are provided in Appendix B.2.3.
@@
In the following section, we elaborate on the taxonomy method for model selection in Section 4, followed by the key design elements of DPImageBench for data preparation, model training, and evaluation in Section 5.
4 Taxonomy and Method Description
This section provides an overview of the taxonomy and elaborates on DP image synthesis algorithms. We present a novel taxonomy of existing methods from two orthogonal dimensions: (1) The Private Data Perspective: where the noise is added, and (2) The Public Data Perspective: how public data is used to enhance DP image synthesis. Note that this paper does not intend to benchmark every possible method but focuses only on state-of-the-art ones. For example, we omitted some GAN- and VAE-based methods, which are not state-of-the-art [4, 18, 31]. Additionally, we include all diffusion-based algorithms due to their exceptional performance. That said, as DPImageBench is a general framework, users can always add more methods to it.
4.1 The Private Data Perspective
We formulate the pipeline of DP image synthesis as follows: public images and sensitive images are fed into the synthesizer for pretraining and finetuning. The well-trained models then generate synthetic images as output. We propose a taxonomy based on where noise is added, as outlined in Table I, categorizing current methods into input-level, model-level, and output-level approaches (Figure 2).
@@ 4.1 The Private Data Perspective
Output-level: PE leverages model inference APIs to generate images and uses each private image to iteratively vote for the most similar API-generated images, forming a voting histogram. Noise is added to this histogram for DP, guiding pretrained models to generate images.
@@
We explain in Appendix E why the state-of-the-art method DP-Promise [24] does not strictly satisfy DP and, therefore, is not included in our analysis.
4.1.1 Input-Level
@@
DP-MERF [28]. This method uses random Fourier feature representations [45] of Gaussian kernel mean embeddings as the high-level statistical characteristics of the image dataset. Kernel mean embedding (KME) [46] maps a probability distribution into a reproducing kernel Hilbert space (RKHS)[47] by computing the mean of kernel evaluations. DP-MERF uses random Fourier features to approximate KME. Under DP constraints, DP-MERF adds noise, determined by the Gaussian mechanism [36], to the KME of sensitive images. Then, DP-MERF trains a generator that minimizes the Maximum Mean Discrepancy (MMD) [48] between the noisy KME and the KME of the generated images.
DP-NTK [29]. This method claimed that directly computing the KME from the original image dataset is challenging due to the high dimensionality of the images [29, 49]. They propose the use of Neural Tangent Kernels (NTKs) [50], which provide an improvement over the Gaussian kernel used in DP-MERF [28]. In particular, DP-NTK introduces the NTK to represent images, whereby the gradient of the neural network function acts as the feature map to extract perceptual features from the original images. Computing KME based on these perceptual features extracted from NTKs yields improved synthetic image quality.
4.1.2 Model-Level
@@
DP-Kernel [17]. DP-MERF matches the KME across all data, which can lead to mode collapse [51], resulting in synthetic images that are overly similar. Besides, using random Fourier features to approximate KME may also cause potential errors. Instead of adding Gaussian noise to the KME of sensitive images in DP-MERF, DP-Kernel proposes directly adding noise to the MMD loss function that measures the discrepancy between the distributions of sensitive and synthetic images in an RKHS to avoid the mentioned issues.
where and represent the distribution of real and synthetic images by the generator.
@@
DP-GAN [30]. As described in Section 2.2, the discriminator in GANs guides the generator to generate images similar to the real images. DP-GAN trains the discriminator network on the sensitive images using DP-SGD, ensuring that the discriminator weights satisfy DP. As presented in Eq. (5), the weight update of the generator solely relies on the discriminator. Due to the post-processing property of DP [52], the generator also satisfies DP.
@@
DPDM [31]. This paper is the first to use DMs [37] as the DP synthesizer. It trains DMs on sensitive images using DP-SGD to ensure DP. Traditional DMs without DP often use small batch sizes and heavily over-parameterized architectures [37, 53]. Under DP, these factors can result in high levels of noise [35], which affects the performance of synthesis. DPDM introduces modifications to DP-SGD specifically tailored for the DP training of DMs to alleviate this issue. In particular, DPDM introduces the noise multiplicity, where the image is re-used for training to learn to denoise at multiple steps. DPDM shows that this modification does not incur additional privacy cost, while reducing the variance of the diffusion training objective.
@@
GS-WGAN [15]. Unlike DPGAN, which applies perturbations to the entire discriminator, GS-WGAN perturbs only the feedback from the discriminator that guides the generator’s training, thereby ensuring DP in synthesizing images. Specifically, GS-WGAN decomposes the generator’s gradient into two components using the chain rule: an upstream gradient (the output of discriminator with respect to the generated image) and a local gradient (the generated image with respect to the generator’s parameters). The upstream gradient is perturbed using GM to satisfy DP, as only the discriminator has direct access to sensitive images during training. By building a “privacy barrier” between the discriminator and the generator, this approach ensures that private information does not leak into the generator. This property allows GS-WGAN to pretrain the discriminators of GAN on sensitive images without considering privacy to warm up the synthesizing.
@@
PDP-Diffusion [32]. This paper first proposes the paradigm of pretraining the DP image synthesizer on a public dataset without privacy concerns. By leveraging this public dataset, the synthesizer can benefit from a broader knowledge base, which may enhance its performance during the subsequent DP-SGD [35] finetuning process on the sensitive dataset. Later methods using DMs, such as DP-LDM [18] and PrivImage [4], are all built upon this paradigm.
@@
DP-LDM [18]. DP-LDM [18] proposes using LDMs as synthesizers instead of standard DMs used in previous works [32]. It further improves efficiency by finetuning only specific components of the pretrained LDM, namely the label embedding module [38] and the attention module [54]. DP-LDM reduces the number of trainable parameters and requires less noise to maintain the same privacy budget.
@@
DP-LoRA [19]. Built on DP-LDM, DP-LoRA uses Low-Rank Adapters (LoRA) [55], a parameter-efficient finetuning approach, to train LDMs, thus reducing the size of trainable parameters and improving the privacy-utility trade-off.
PrivImage [4]. This paper highlights that a dataset with a distribution similar to the sensitive dataset is more suitable for pertaining [11, 13]. This insight motivates PrivImage to select pretraining data with a distribution that closely matches the sensitive dataset. PrivImage proposes to describe the distribution of image datasets using semantics [56]. Unlike focusing on the low-level features such as color, texture, and edges that can be directly extracted from an image, semantics provides a high-level understanding of images, including their “meaning” or “content” [56].
To access the semantic distribution of images, PrivImage first establishes a semantic query function using the public dataset . This function can be an image captioning method [57, 58] or an image classifier. The objective of query function is defined as [4]:
Here, represents the public dataset, where each entry consists of an image and its associated semantic labels . denotes the number of possible semantic labels in . A label indicates that the image possesses the -th semantic label, while signifies its absence. The function represents the probability that image has semantic . PrivImage extracts semantics from sensitive images using a query function, adds Gaussian noise to the queried semantic distribution to ensure DP, and then selects a subset of public images for pretraining based on this noisy distribution. Finally, PrivImage pretrains the synthesizer on the selected dataset and then fine-tunes the model on the sensitive dataset using DP-SGD [35].
4.1.3 Output-Level
These papers introduce noise during the synthesis process to assist in selecting synthetic images that are most similar to the sensitive images under DP.
@@ 4.1.3 Output-Level
Private Evolution (PE) [5]. This method uses the foundation APIs, which can come from either blackbox API-based models such as DALLE-2 [41] or open-sourced models like Stable Diffusion [38]. PE iteratively uses each private image to vote for the most similar images generated by the APIs, forming a voting histogram. DP is ensured by adding noise to this histogram. Using the noisy histogram, PE identifies the images most similar to the private ones and then prompts the pretrained models to generate more variations of them. As a result, PE can be applied to any generative model capable of producing variations of an image (such as diffusion models by adding noise to the image and using the model to denoise it). Notably, PE relies solely on model inference and does not require any training on sensitive images. Lin et al. [5] prove that the distribution of images generated by PE can converge to that of the private images.
4.1.4 Mix-Level
@@
DP-FETA [26]. This method enables synthesizers to learn image features progressively from simple to complex. DP-FETA divides the training process for generating complex images into two stages: (1) DP-FETA extracts basic features, such as rough outlines and colors, from sensitive images using “central images”—representations of central tendency measures (e.g., mean, mode). These images, perturbed with Gaussian noise at the “input level” to ensure DP, build a shortcut to facilitate early-phase synthesizer training. After the training on central images, the diffusion models can generate rough and statistically imperfect images. (2) The synthesizer is then fine-tuned on original sensitive images using DP-SGD [35] to generate more realistic images, with noise added at the “model-level.”
@@
4.2 The Public Data Perspective
As shown in Table II, pretraining synthesizers on public datasets, such as ImageNet [59], which are openly available and do not raise privacy concerns, can enhance the utility and fidelity of DP synthetic images [4, 32, 18]. We believe that pretraining brings new insight into the community. Previous taxonomy lacks a systematic analysis of how public data is utilized [20, 1]. We fill this gap and outline potential directions as follows.
Pretrained model: PE [5] utilizes inference APIs of foundation models, which have been trained on public datasets.
@@ 4.2 The Public Data Perspective
Public dataset with selection: PrivImage [4] shows that selecting a subset of pretraining images whose distribution closely matches the sensitive dataset yields better performance compared to using the entire public dataset.
This taxonomy provides insights for developing new technologies. (1) Inspired by the public dataset selection approach in PrivImage, one potential avenue is to refine methods that do not consider selecting pretrained models (like PE). (2) Integrating the public dataset selection into approaches that have not yet been considered could be another promising direction. Specifically, in Section 7.3, we investigate how incorporating the public dataset selection approach from PrivImage into methods without selection can enhance synthetic performance.
5 Key Designs of DPImageBench
5.1 Data Preparation
DPImageBench offers a flexible tool that allows users to select both public and sensitive datasets. This setup facilitates the pretraining of DP image synthesizers on public datasets for all studied algorithms, including those that were not initially designed to leverage public datasets for pretraining, e.g., DP-MERF [28], DP-NTK [29], etc. The README.md file of our repository provides instructions for users to customize their dataset.
5.2 Model Training
5.2.1 The Failure Probability
For a meaningful DP guarantee, should satisfy that , where presents the size of the sensitive dataset [60]. Many previous works ignore this relationship and use . In DPImageBench, we set following prior works [60, 9].
@@ 5.2.1 The Failure Probability
DP-MERF [28] | ✗ | Acc | ✗ | |
DP-NTK [29] | ✗ | Acc | ✗ | |
DP-Kernel [17] | FID | Acc | ✗ | |
PE [5] | FID | Acc | Nearest Sample | |
GS-WGAN [15] | FID, IS | Acc | ✗ | |
DP-GAN [30] | FID | Acc | ✗ | |
DPDM [31] | FID | Acc | ✗ | |
DP-FETA [26] | FID | Acc | ✗ | |
PDP-Diffusion [32] | FID | Acc | ✗ | |
DP-LDM [18] | FID | Acc | ✗ | |
DP-LoRA [19] | FID | Acc | ✗ | |
DP-Promise [24] | FID, IS | Acc | ✗ | |
PrivImage [4] | FID | Acc | MIA |
5.2.2 Unified Framework
Failing to standardize the pretraining dataset and model architectures can result in unfair comparisons. Even when the same model type is used as the synthesizer, variations in the hyper-parameters of the synthesizer and DP-SGD (e.g., batch sizes, epochs, and learning rates) can greatly influence performance. DPImageBench standardizes the public images for pretraining and model architectures to ensure fair and consistent comparisons.
In DP-LDM and DP-LoRA, the synthesizer uses stable diffusion [38]. In contrast, other methods, such as DPDM [31] and PrivImage [4], rely on standard DMs. As described in Section 2.2, LDMs are implemented by incorporating an autoencoder into DMs. Therefore, DPImageBench provides standardized implementations for the common components of both LDMs and standard DMs, including the structure of the noise prediction network and hyper-parameters of DP-SGD. We also implement DP-LDM with standard DMs to examine the extent to which performance improvements are influenced by differences in the synthesizer architecture. We refer to the DP-LDM that applies standard DMs as “DP-LDM (SD).”
5.2.3 Extensions to New Algorithms

### v3

@@ 3. Overview of DPImageBench
DPImageBench should comprehensively evaluate synthetic images generated by different algorithms for comparative studies.
@@
(1) Model Selection. DPImageBench allows users to select among twelve distinct DP image synthesis algorithms, which are categorized into three groups as introduced in Section 4. DPImageBench also provides simple interfaces to integrate new algorithms.
(2) Data Preparation. This module prepares the public images and sensitive images for the following modules by partitioning and preprocessing the data, such as normalization. DPImageBench allows users to select public and sensitive datasets flexibly.
(3) Model Training. This module involves (optional) synthesizer pretraining on the public dataset and the main training on the sensitive dataset. We unify the model architectures and hyper-parameters across different methods. Besides, we provide flexible configuration files, allowing users to adjust hyper-parameters according to their preferences. DPImageBench supports pretraining for all algorithms, even those not originally designed with pretraining. Details are provided in Appendix B.2.3.
@@
4. Taxonomy and Method Description
We present a novel taxonomy of existing methods from two orthogonal dimensions: (1) The Private Data Perspective: where the noise is added, and (2) The Public Data Perspective: how public data is used to enhance DP image synthesis. Note that this paper does not intend to benchmark every possible method but focuses only on state-of-the-art ones. For example, we omitted some GAN- and VAE-based methods, which are not state-of-the-art (Li et al., 2024; Liu et al., 2023; Dockhorn et al., 2023). , and we include all diffusion-based algorithms due to their exceptional performance. That said, as DPImageBench is a general framework, users can always add more methods to it.
4.1. The Private Data Perspective
We formulate the pipeline of DP image synthesis as follows: public images and sensitive images are fed into the synthesizer for pretraining and finetuning. The well-trained models then generate synthetic images as output. We propose a taxonomy based on where noise is added, categorizing current methods into input-level, model-level, and output-level approaches (Figure 2).
@@ 4.1. The Private Data Perspective
Output-level: PE leverages model inference APIs to generate images, iteratively voting for the most similar API-generated images per private image to create a histogram. Noise is added to this histogram for DP, guiding pretrained models to produce images.
@@
We explain in Appendix E why DP-Promise (Wang et al., 2024) does not strictly satisfy DP and, therefore, is not included in our analysis.
4.1.1. Input-Level
@@
DP-MERF (Harder et al., 2021). This method uses random Fourier feature representations (Rahimi and Recht, 2007) of Gaussian kernel mean embeddings as the high-level statistical characteristics of the image dataset. Kernel mean embedding (KME) (Smola et al., 2007) maps a probability distribution into a reproducing kernel Hilbert space (RKHS) (Aronszajn, 1950) by computing the mean of kernel evaluations. DP-MERF uses random Fourier features to approximate KME. DP-MERF adds noise, determined by the Gaussian mechanism (Mironov et al., 2019), to the KME of sensitive images. Then, DP-MERF trains a generator that minimizes the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) between the noisy KME and KME of generated images.
DP-NTK (Yang et al., 2023). This method claimed that computing the KME directly from high-dimensional image datasets is challenging (Yang et al., 2023; Harder et al., 2023). It proposes the use of Neural Tangent Kernels (NTKs) (Arora et al., 2019), outperforming the Gaussian kernel in DP-MERF (Harder et al., 2021). DP-NTK uses NTKs to represent images, whereby the gradient of the neural network function acts as the feature map to extract perceptual features from the original images. Computing KME from NTKs yields improved synthetic image quality.
4.1.2. Model-Level
@@
DP-Kernel (Jiang et al., 2023). DP-MERF matches the KME across all data, which can lead to mode collapse (Srivastava et al., 2017), resulting in synthetic images that are overly similar. Besides, using random Fourier features to approximate KME may also cause potential errors. Instead of adding Gaussian noise to the KME of sensitive images, DP-Kernel proposes directly adding noise to the MMD loss function that measures the discrepancy between the distributions of sensitive and synthetic images in an RKHS to avoid the mentioned issues.
where and mean the distribution of real and synthetic images.
@@
DP-GAN (Xie et al., 2018). As described in Section 2.2, the discriminator in GANs guides the generator to generate images similar to the real images. DP-GAN trains the discriminator network on the sensitive images using DP-SGD, ensuring that the discriminator weights satisfy DP. As presented in Eq. (5), the weight update of the generator solely relies on the discriminator. Due to the post-processing property of DP (Dwork et al., 2014), the generator also satisfies DP.
@@
DPDM (Dockhorn et al., 2023). This paper pioneers DMs (Ho et al., 2020) as DP synthesizers. It trains DMs on sensitive images using DP-SGD to ensure DP. Traditional DMs, without DP, use small batches and over-parameterized models (Ho et al., 2020; Karras et al., 2022), but under DP, this leads to noisy gradients (Abadi et al., 2016), degrading synthesis. DPDM introduces the noise multiplicity, where the image is re-used for training to learn to denoise at multiple steps. DPDM shows that this modification does not incur additional privacy costs while reducing the variance of the gradient.
@@
GS-WGAN (Chen et al., 2020). Unlike DPGAN, which perturbs the entire discriminator, GS-WGAN perturbs only the discriminator’s feedback guiding the generator. It splits the generator’s gradient via the chain rule into an upstream gradient (the output of discriminator with respect to the generated image) and a local gradient (the generated image with respect to the generator’s parameters), perturbing the upstream gradient with GM for DP, as only the discriminator accesses sensitive images. This “privacy barrier” prevents leakage to the generator, enabling pretraining of discriminators on sensitive images without privacy concerns to warm up the synthesizing.
@@
PDP-Diffusion (Ghalebikesabi et al., 2023). This paper first proposes the paradigm of pretraining the DP image synthesizer on a public dataset without privacy concerns. By leveraging this public dataset, the synthesizer can benefit from a broader knowledge base, which may enhance its performance during the subsequent DP-SGD (Abadi et al., 2016) finetuning process on the sensitive dataset. Later methods using DMs, such as DP-LDM (Liu et al., 2023) and PrivImage (Li et al., 2024), are all built upon this paradigm.
@@
DP-LDM (Liu et al., 2023). DP-LDM (Liu et al., 2023) proposes using LDMs as synthesizers instead of standard DMs. It further improves efficiency by finetuning only specific components of the pretrained LDM, namely the label embedding module (Rombach et al., 2022) and the attention module (Vaswani, 2017). DP-LDM reduces the number of trainable parameters and requires less noise to maintain the same privacy budget.
@@
DP-LoRA (Tsai et al., 2024). Built on DP-LDM, DP-LoRA leverages Low-Rank Adapters (LoRA) (Hu et al., 2022), a parameter-efficient finetuning approach, to train LDMs, thus reducing the size of trainable parameters and improving the privacy-utility trade-off.
This paper shows that pretraining on a dataset with a distribution similar to the sensitive dataset is more effective (Yu et al., 2023; Yin et al., 2022). PrivImage thus selects pretraining data matching the sensitive dataset’s distribution, using semantics (Long and Lu, 2021) to capture high-level image meaning rather than low-level features like color or texture. PrivImage defines a semantic query function on public dataset , using image captioning (Vinyals et al., 2015; Xu et al., 2015) or classification, with the objective:
Here, is an image, means semantic labels ( if the -th label applies, else ), and is the probability of label . PrivImage queries semantics from sensitive images; adds Gaussian noise for DP; selects public images based on noisy distributions; and fine-tunes the pretrained model on sensitive datasets using DP-SGD (Abadi et al., 2016).
4.1.3. Output-Level
This kind of method introduces noise during the synthesis process under DP.
@@ 4.1.3. Output-Level
Private Evolution (PE) (Lin et al., 2024). This method uses foundation APIs from either black-box models like DALL-E 2 (Ramesh et al., 2022) or open-source models such as Stable Diffusion (Rombach et al., 2022). PE iteratively employs each private image to vote for the most similar API-generated images, constructing a voting histogram. DP is achieved by adding noise to this histogram. Using the noisy histogram, PE identifies images closely resembling the private ones and prompts pretrained models to generate additional variations. Consequently, PE is adaptable to any generative model capable of producing image variants (e.g., diffusion models that denoise noisy images). Notably, PE operates solely via model inference, requiring no training on sensitive images. Lin et al. (Lin et al., 2024) prove that the distribution of images generated by PE can converge to that of the private images.
4.1.4. Mix-Level
@@
DP-FETA (Li et al., 2025). This method allows synthesizers to learn image features from simple to complex. DP-FETA splits training into two stages: (1) It extracts basic features (e.g., outlines, colors) from sensitive images using “central images”—representations of central tendency measures (e.g., mean, mode). These images are perturbed with Gaussian noise at the input level for DP, building a shortcut to facilitate early-phase synthesizer training. After the training on central images, the diffusion models can generate rough and statistically imperfect images. (2) Fine-tuning on sensitive images with DP-SGD (Abadi et al., 2016) adds model-level noise to produce realistic images.
@@
4.2. The Public Data Perspective
As shown in Table 1, pretraining synthesizers on public datasets, such as ImageNet (Deng et al., 2009), which are openly available and do not raise privacy concerns, can enhance the utility and fidelity of DP synthetic images (Li et al., 2024; Ghalebikesabi et al., 2023; Liu et al., 2023). We believe that pretraining brings new insight into the community. Previous taxonomy lacks a systematic analysis of how public data is utilized (Chen et al., 2023; Hu et al., 2024). We fill this gap and outline potential directions as follows.
Pretrained model: PE (Lin et al., 2024) utilizes inference APIs of foundation models, which have been trained on public datasets.
@@ 4.2. The Public Data Perspective
Public dataset with selection: PrivImage (Li et al., 2024) shows that selecting a subset of pretraining images whose distribution closely matches the sensitive dataset yields better performance compared to using the entire public dataset.
This taxonomy provides insights for developing new technologies. (1) Inspired by the public dataset selection approach in PrivImage, one potential avenue is to refine methods that do not consider selecting pretrained models (like PE). (2) Integrating the public dataset selection into approaches that have not yet been considered could be another promising direction. Section 7.3 investigates how incorporating the public dataset selection from PrivImage into methods without selection can enhance synthetic performance.
5. Key Designs of DPImageBench
5.1. Data Preparation
DPImageBench offers a flexible tool that allows users to select both public and sensitive datasets. This setup facilitates the pretraining of DP image synthesizers on public datasets for all studied algorithms, including those that were not initially designed to leverage public datasets for pretraining, e.g., DP-MERF (Harder et al., 2021), DP-NTK (Yang et al., 2023), etc. The README.md file of our repository provides instructions for users to customize their dataset.
5.2. Model Training
5.2.1. The Failure Probability
For a meaningful DP guarantee, should satisfy that , where presents the size of the sensitive dataset (Lin et al., 2021a). Many previous works ignore this relationship and use . In DPImageBench, we set following prior works (Lin et al., 2021a; Yue et al., 2023).
@@ 5.2.1. The Failure Probability
| DP-MERF (Harder et al., 2021) | ✗ | Acc | ✗ | |
| DP-NTK (Yang et al., 2023) | ✗ | Acc | ✗ | |
| DP-Kernel (Jiang et al., 2023) | FID | Acc | ✗ | |
| PE (Lin et al., 2024) | FID | Acc | Nearest Sample | |
| GS-WGAN (Chen et al., 2020) | FID, IS | Acc | ✗ | |
| DP-GAN (Xie et al., 2018) | FID | Acc | ✗ | |
| DPDM (Dockhorn et al., 2023) | FID | Acc | ✗ | |
| DP-FETA (Li et al., 2025) | FID | Acc | ✗ | |
| PDP-Diffusion (Ghalebikesabi et al., 2023) | FID | Acc | ✗ | |
| DP-LDM (Liu et al., 2023) | FID | Acc | ✗ | |
| DP-LoRA (Tsai et al., 2024) | FID | Acc | ✗ | |
| DP-Promise (Wang et al., 2024) | FID, IS | Acc | ✗ | |
| PrivImage (Li et al., 2024) | FID | Acc | MIA |
5.2.2. Unified Framework
Inconsistent pretraining datasets and model architectures can lead to unfair comparisons. Even with identical synthesizer types, differences in hyperparameters for the synthesizer and DP-SGD—such as batch sizes, epochs, and learning rates—can significantly affect performance. DPImageBench standardizes public pretraining images and model architectures to ensure fair evaluations.
In DP-LDM and DP-LoRA, the synthesizer uses Stable Diffusion (Rombach et al., 2022), whereas methods like DPDM (Dockhorn et al., 2023) and PrivImage (Li et al., 2024) utilize standard diffusion models (DMs). As outlined in Section 2.2, LDMs integrate an autoencoder into DMs. To facilitate fair comparisons, DPImageBench provides standardized implementations for shared components of LDMs and standard DMs, including the noise prediction network structure and DP-SGD hyperparameters. Additionally, we implement DP-LDM with standard DMs—termed “DP-LDM (SD)”—to assess how synthesizer architecture variations impact performance improvements.
5.2.3. Extensions to New Algorithms